Circular No. 6427

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions)

BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science) Phone 617-495-7244/7440/7444 (for emergency use only)

GRS 1739–278 I. F. Mirabel and J. Marti, Service d’Astrophysique, Centre d’Etudes

de Saclay; P. A. Duc, European Southern Observatory (ESO); and L. F. Rodriguez, Universidad Nacional Autonoma de México, communicate: “Deep images in the bands I, J, and K were obtained with several ESO telescopes in the context of a TOO program in a search for optical and/or infrared counterparts of the x-ray/radio nova GRS 1739–278 (IAUC 6348, 6383). The first observations were on Apr. 1, one day after the ROSAT detection of a bright x-ray source (IAUC 6426), and they were continued for an interval of several weeks. We find a possible highly-reddened counterpart in all observed bands within 1′′ of the radio counterpart at α = 17^{h}42^{m}40^{s}.0, δ = −27^{o}44′53′′ (equinox 2000.0). Mean approximate magnitudes are R = 20.5, J = 16.0, K = 14.7. An optical spectrum obtained on Apr. 24–25 with the 3.6-m telescope by H. Duerbeck and S. Benetti showed no strong Hα emision line.”

SUPERNOVA 1996ah IN NGC 5640 G. V. Williams, Harvard-Smithsonian Center for Astrophysics, pro-

vides his measurement of the position of SN 1996ah from a CCD image taken by P. Challis on Apr. 16.127 UT with the 1.2-m reflector at Mt. Hopkins: α = 14^{h}20^{m}39^{s}.02, δ = +80^{o}07′21′′.0 (equinox 2000.0), which is 5′′.0 west and 1′′.5 south of the galaxy’s center.

D. D. Balam, University of Victoria, reports his measurement of the position (end figures) for SN 1996ah from images obtained by G. C. L. Aikman and F. Dossa with the 1.82-m Plaskett telescope (Dominion Astrophysical Observatory) on June 25.38 UT: 38^{s}.83, 21′′.0 (which is 5′′.7 west and 1′′.4 south of the galaxy’s center).

SUPERNOVA 1996ai IN NGC 5005 S. Pesci, Milan, reports that P. Mazza observed NGC 5005 on June

13.9 UT with a 0.40-m telescope and saw nothing at the position of SN 1996ai to a limiting visual magnitude of 15.0 (poor sky conditions).

COMET 78P/GEHRELS 2 CCD magnitude estimates: June 18.37 UT, 20.4 (W. Offutt, Cloud-

croft, NM, 0.6-m f/7 reflector); 26.34, 21.0 (C. W. Hergenrother, Mt. Hopkins, AZ, 1.2-m reflector; stellar appearance).

1996 July 1  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1996 CBAT  Daniel W. E. Green